# System 0/1/2/3: Quad-process theory for multi-timescale embodied collective cognitive systems

### v1

@@ 2.3 Multi-agent Systems with Emergent Language
Another promising direction involves fostering collaboration through debates or discussions among agents to improve their reasoning capabilities. Chen et al., [19] proposed RECONCILE, a multi-model, multi-agent framework designed as a roundtable discussion among diverse LLM agents, significantly enhancing their reasoning capabilities. Smith and Doe, [121] introduced Multi-Agent LLM Training (MALT), a training approach where multiple agents collaborate in a debate setting to solve reasoning tasks. Inspired by neural scaling laws, Qian et al., [108] explored how increasing the number of agents in multi-agent collaboration impacts performance. In addition, Subramaniam et al., [126] proposed the use of social interactions in multi-agent debates to fine-tune LLMs.
@@ 2.4 Bergson’s Philosophy: Multi-timescale interpretation
The temporal hierarchy related to life and mind is an important topic in philosophy. French philosopher Henri Bergson (1859–1941) is renowned for his distinctive theory of time. Recent studies, especially from Project Bergson in Japan (PBJ), have provided new insights into his work ([55], [92], [135], [36], [161], [52], [53]). This new interdisciplinary trend is called "Expanded Bergsonism" ([91]). According to Hirai ([52]), Bergson’s philosophy of time differs from traditional approaches by endorsing multiple intrinsic timescales. This subsection outlines Bergson’s "Multi-Timescale" (MTS) interpretation in relation to System 0/1/2/3.
@@
Bergson’s concept of durée (duration), initially introduced to describe the human conscious experience of time, was expanded in his second major book, Matter and Memory ([9]), to encompass the temporality of matter and different life forms from the perspectives of emergence and evolution. This idea of ““multiple rhythms of durées”” can be viewed today as a pluralism of timescales.
@@
In traditional philosophy, time has been treated as either a single cosmic time (absolute time) or psychological time experienced subjectively, often posited in dualistic opposition. The notion of pluralistic intrinsic timescales has been largely ignored. In contrast, Bergson focused on proper timescales inherent to each local interaction system. His famous reference to "sugar dissolving in water" illustrates that the event exhibits a duration "which cannot be extended or shortened at will" ([10], p. 10). Different solutes and solvents require different times. Bergson’s philosophy aligns well with the concept of "characteristic time" in thermodynamics and condensed matter physics crucial for understanding system responses and relaxation processes.
@@
Importantly, cognitive activities are natural phenomena with intrinsic timescales. Just as heat diffusion or chemical reactions require specific durations, organisms require specific intervals to perceive situations and perform actions. Thus, sensorimotor processes can be understood as ““relaxation times”” required to re-stabilize the system. Bergson defines the ““thickness”” of the present this way: the real present ““necessarily occupies a duration”” and it is ““in its essence, sensorimotor”” ([9], p. 153). While he does not specify numbers, contemporaries like Wilhelm Wundt and modern findings suggest that the "psychological present" spans approximately 3-5 s ([102], [30], [155]).
@@
Temporal consciousness (durée) resides within this thick present. Changes too slow, such as celestial rotation or plant growth, are not directly experienced as "time perception." Conversely, there is also a lower limit: the minimal temporal size of perceptual elements. Bergson referred to "temporal resolution," citing a 2-ms value ([9], p. 231, [29]). Rapid changes, as seen in fluorescent lights, also escape our time perception, leading to phenomena such as the wagon-wheel effect.
@@
Thus, cognition operates within a specific temporal domain defined by upper and lower limits, termed here as the "window of the present." Consciousness perceives temporal changes only within this specific domain. Bergson views this timescale-locality not as a constraint but as an enabling condition for consciousness. It is this specific "slowness" relative to microscopic physical phenomena that is key to the very emergence of consciousness([51], [52]).
@@
Besides, as Hirai ([54]) clarified, Bergson had already discussed a dual process of cognition based on the flexible nature of the window of the present: "automatic recognition," which remains confined to the thin present, and "attentive recognition," which leverages the thickness of the present (Bergson, 1896, chap. 2), approximately corresponding to modern Systems 1 and 2, respectively. This point will be revisited in Section 4.2.
@@
Bergson’s philosophy of time, particularly its emphasis on multi-timescale temporality, aligns with aspects of dual process theory. However, it also offers the potential for broader discussions concerning the philosophy of consciousness and more expansive concepts of time. Expanding dual process theory while elucidating its connections with Bergson’s philosophy of time could prove crucial for advancing philosophical discourse in these areas.
@@ 3.2 Collective Predictive Coding and System 3
The subject of System 3 shifts the perspective from individual life to life as a collective entity. This can be understood by considering the question “Who is minimizing the free energy* described in Equations 5-9?” The entity minimizing free energy is not the individual but rather the collective agents who share the variable $w$ defined in Equation 5, representing a collective form of life that engages in predictive coding of the world. While System 3’s operation corresponds to the dynamics of symbol and language emergence, one might assume that its learning relies solely on System 2 processes, such as individual language learning. However, we assert that System 3 is a collective emergent system. It adapts to its environment not only through the internal neural plasticity of individual agents but also by leveraging the plasticity of language itself in predictive coding. Consequently, the language formed through this process reflects collective predictive coding as a group.
@@ 3.3 Relationships and Dynamics Among Systems: A Cognitive and Temporal Perspective
4 Extended Bergson’s MTS for System 0/1/2/3
Bergson’s philosophy provides a grounding framework for the above-discussed extended four-tiered system in philosophical terms. This is because his philosophy of time encompasses not only the "fast and slow" in conscious cognition, which corresponds to his second major work Matter and Memory, but also questions of evolution and society, corresponding to his third and fourth major works, respectively. In this section, we introduce Bergson”s approach, which combines the ““landscape,”” an extended notion of space, with the ““window of the present,”” rendered elastic, as a means to structure and clarify the conceptual foundations behind such a wide range of phenomena.
@@ 4.1 Three Axes of Timescales
Bergson’s philosophy of time offers a broad perspective on various phenomena, ranging from matter to society, at different timescales. When distinguishing between different timescales, there are multiple axes that should not be conflated, and differentiating these is beneficial for our discussion (see Figure 5).
@@
The first series refers to the multiple temporal layers that structure a biological individual. Centered around the present window, which is framed by the aforementioned layers 1 and 2, it comprises layer 0, a material layer below, and layer 3, a memory layer above (see the MTS structure on the right of Figure 5). The second series represents the timescales required for the formation of adaptive optimization (““landscape”” formation or ““canalization,”” as discussed later; see left of Figure 5). For example, evolution involves timescales ranging from thousands to millions of years, while learning involves timescales ranging from weeks to months. The third series represents the present window size (middle of Figure 5). A thicker present window allows for real-time access to a richer reservoir of pure memory in layer 3. Strictly speaking, this is not a difference in time "scales," but it serves as a decisive factor influencing higher-order emergent properties, such as deliberation and verbal ingenuity.
In real-world phenomena, multiple timescales from these different series are intertwined and manifested together. Using the systems proposed in this paper as examples, the following applies: Both Systems 0 and 1 utilize only a "thin" window of the present, but the timescales of the underlying canalization differ between evolution and learning. System 2 has a "thick" present window, enabling access to memory, and interacts with the learning landscape. System 3 has an "extended" present window and additionally interacts with a collectively formed societal landscape.
@@ 4.2 Two Types of Memory and Thickness of the Present
Bergson’s theory of time explains natural phenomena through the combined effect of two types of "memory." This distinction is often referred to as a precursor to modern procedural and episodic memory ([146], [106]); however, it has a broader scope. One type is "motor memory," which refers to a series of motor modules learned through repetition, such as skills and habits in humans. The other is "pure memory," which preserves original experiences before they are organized into such patterns. For example, through instances of bicycle riding trials (r1, r2, r3), a motor memory for riding a bicycle (R) is acquired. However, one can also recall r1 as a specific event in life on a particular date. Pure memory is what retains these instances.
The critical point is that the organization of motor memory, as an adaptation to the environment, operates under the constraints of utility at the given moment, whereas pure memory provides a temporally open resource, offering a surplus beyond these established typical patterns. This surplus allows for flexibility and adaptability in novel situations that may be encountered later.
Let us now revisit the two types of recognition briefly mentioned in Section 2.4: a fast but limited automatic process and a slow but intentional process, the latter being referred to as "attentive recognition" ([9], chap. 2). Both automatic and attentive recognition are based on motor memory and share the characteristic of preparing interactions with stereotyped environments. However, automatic recognition, which operates with a "thin" window of the present, is confined to immediate motor responses. In contrast, attentive recognition, with its "thick" window of the present, allows access to pure memory. This temporal opening enables the active reconstruction of the recognized object as a mental image and, in some cases, even brings the individual back to a scene from their personal history. Such a temporal characterization would also likely apply to Systems 1 and 2.
Motor memory is essential for smoothing our immediate cognition and actions. Without this, complex motor responses would have to be organized from scratch each time, making hierarchical and intricate processes unattainable. However, if all experiences were solely dedicated to building motor memory, leaving no surplus, there would be no additional resources to draw upon in new situations beyond its scope. The size of the present window acts as a constraint on this real-time improvisation. These dual-memory resources are extremely important for our intellectual pursuits and creativity. They allow us to be prepared and supported by motor memory while retaining the freedom to depart from it when necessary.
@@ 4.3 Extended Definition of Motor Memory: Canalized Landscape
In this paper, we define motor memory as self-organized "prior optimization" that occurs before actual experience (here, optimization does not necessarily aim for the global optimum). Mastering a movement and learning its patterns reduce the cognitive and motor control load when a situation arises. The effect of motor memory is clear: temporal outsourcing of optimization. This suggests that motor memory is not limited to internal mechanisms, such as updating motor programs or world models within an organism. Organisms, by modifying their environment through behaviors such as marking territory or building nests—and in humans, through experiences such as roads, buildings, and product design—create conditions that allow for automatic responses without requiring conscious decisions.
@@
This definition, which focuses on the underlying temporal structure rather than the phenomenological classification of events, can also enable us to capture adaptive optimization on an evolutionary scale as a large-scale self-organizing landscape. This leads to unconscious behaviors based on innate behavioral and morphological traits, which, in terms of this paper, correspond to System 0 (Bergson calls this ““instinct””).
@@
Bergson cited the Sphex wasp stinging a caterpillar’s nine nerve centers in a specific order to paralyze but not kill it ([10], pp. 174–175). This intricate maneuver is not learned or reasoned but pre-configured. On an evolutionary scale, this has been achieved through countless variations and selections. Behavior optimization here is completed even prior to the organism”s birth.
@@
Thus, at both the individual and evolutionary levels, establishing stable sensorimotor routes in advance within the environment eliminates the need for the present window. In this sense, Bergson argues that as motor memory becomes perfected, it "steps outside of time."([9], p. 91) To illustrate this point, we introduce the concept of the ““Canalized Landscape.””
@@
Bergson’s theory of perception involves the idea of knowledge outsourcing. In Matter and Memory, he rejects both representationalism and atomistic associationism, developing a form of "direct realism" grounded in the possible actions of organisms. The notion that potential behavioral inducements are structured within the environment not only anticipates von Uexküll’s concept of Umwelt ([149]) but also aligns remarkably well with the dynamical systems perspective discussed in Section 2.1. Each organism operates within its own action-inducing space, extended beyond the body, and is shaped through motor memory, such as evolution and learning. Bergson likened its formation process to river formation over geological timescales, calling it canalization ([9], p. 280; [10], pp. 94-95). Following C. H. Waddington, a renowned geneticist who inherited this idea via A. N. Whitehead ([107]), we term the resulting space the "landscape" ([150]).
@@
Thanks to this canalized landscape, an organism does not need to build its behavior from scratch in each instance. Instead, organisms pre-construct an appropriate landscape in which the potential for appropriate actions is well arranged. Canalization prepares this potential space, and physical laws, evolutionary adaptations, and learning all contribute to it. Ideally, organisms achieve locally optimized behaviors by simply "rolling down" the landscape.
@@
If a river meanders suitably, water does not need to "think"; by flowing towards lower potential, it naturally follows the optimized ““valley”” paths. Similarly, the Sphex wasp’s behavior is not solely due to internal motor control but also its coupling with the caterpillar’s body shape and structure, allowing behavior to flow automatically. Here, the spatial distinction between the internal and external is less important; the focus is temporally outsourcing landscape formation (canalization) prior to the present window.

### v2

@@ 2.3 Multi-agent Systems with Emergent Language
Another promising direction involves fostering collaboration through debates or discussions among agents to improve their reasoning capabilities. Chen et al., [19] proposed RECONCILE, a multi-model, multi-agent framework designed as a roundtable discussion among diverse LLM agents, significantly enhancing their reasoning capabilities. Motwani et al., [93] introduced Multi-Agent LLM Training (MALT), a training approach where multiple agents collaborate in a debate setting to solve reasoning tasks. Inspired by neural scaling laws, Qian et al., [108] explored how increasing the number of agents in multi-agent collaboration impacts performance. In addition, Subramaniam et al., [125] proposed the use of social interactions in multi-agent debates to fine-tune LLMs.
@@ 2.4 Bergson’s Philosophy: Multi-timescale interpretation
The temporal hierarchy related to life and mind is an important topic within philosophy as well. French philosopher Henri Bergson (1859–1941) is renowned for his distinctive theory of time. Recent studies, particularly from the Project Bergson in Japan (PBJ), have brought new insights into his work [54, 91, 134, 35, 160, 51, 52]. This new interdisciplinary approach is termed “Expanded Bergsonism” [90]. According to Hirai, [51], Bergson’s philosophy of time differs from traditional approaches by emphasizing multiple intrinsic timescales. This subsection outlines Bergson’s Multi-Timescale (MTS) interpretation in relation to System 0/1/2/3. The multi-timescale interpretation inspired by Bergson aligns naturally with contemporary cognitive architectures and robotics frameworks, particularly the hierarchical temporal dynamics observed in neuro-dynamical systems such as MTRNN and RNNPB, discussed in Section 2.1.2.
@@
Bergson’s concept of durée (duration), initially introduced to describe human conscious experience of time, was expanded in his second major book Matter and Memory [9] to encompass the temporality of matter and various life forms from the perspectives of emergence and evolution. This concept of “multiple rhythms of durées” can today be interpreted as a pluralism of timescales.
@@
Traditional philosophical treatments of time have generally viewed it either a single cosmic absolute time or as subjective psychological time, often placing them in dualistic opposition. The notion of multiple intrinsic timescales has largely been overlooked. In contrast, Bergson focused on the proper timescales inherent to each local interaction system. His renowned example of “sugar dissolving in water” illustrates that each event has its own duration, “which cannot be extended or shortened at will” ([10], p. 10). Different solutes or solvents require different times. Bergson’s philosophy resonates well with the concept of “characteristic time” in thermodynamics and condensed matter physics—concepts crucial to understanding system responses and relaxation processes.
@@
Importantly, cognitive activities themselves are natural phenomena that possesses intrinsic timescales. Just as heat diffusion or chemical reactions require a specific durations, organisms also need specific intervals to perceive situations and perform actions. Thus, sensorymotor processes can be viewed as “relaxation times” required to re-stabilize the system. Bergson defined the “thickness” of the present as follows: the real present “necessarily occupies a duration” and is “in its essence, sensorimotor” ([9], p. 153). While he did not specify precise values, contemporary scholars such as Wilhelm Wundt and recent studies suggest the “psychological present” spans approximately 3-5 seconds [102, 30, 154].
@@
Within this thick present, temporal consciousness (durée) resides. Changes that are too slow, such as celestial rotation or plant growth, are not directly experienced as “time perception.” Conversely, there is also a lower limit—the minimal temporal unit of perceptual elements. Bergson refers to “temporal resolution,” citing a value of around 2 millisecond ([9], p. 231, [29]). Changes occurring more rapidly, such as those in fluorescent lighting, similarly escape our perception of time, resulting in phenomena like the wagon-wheel effect.
@@
Thus, cognition operates within a specific temporal domain bounded by these upper and lower limits, termed here as the “window of the present.” Our consciousness perceives temporal changes exclusively within this domain. Bergson viewed this timescale-locality not as a limitation but as an enabling condition for consciousness itself. This particular “slowness,” relative to microscopic physical phenomena, is critical for the very emergence of consciousness [50, 51].
@@
Moreover, as Hirai ([53]) clarified, Bergson also described a dual-process model of cognition based on the flexible nature of the window of the present: “automatic recognition,” which remains confined to the thin present, and “attentive recognition,” which leverages the thickness of the present, roughly corresponding to modern Systems 1 and 2 (Bergson, 1896, chap. 2). This point will be revisited in Section 4.2.
@@
Bergson’s philosophy of time, particularly with its emphasis on multi intrinsic timescales, aligns with certain aspects of dual-process theories. Furthermore, it offers potential for broader discussions concerning the philosophy of consciousness and more expansive conceptions of time. Extending dual-process theory and elucidating its connections to Bergson’s philosophy of time could prove crucial for advancing philosophical discourse in these areas.
@@ 3.2 Collective Predictive Coding and System 3
The subject of System 3 shifts the perspective from individual life to life as a collective entity. This can be understood by considering the question “Who is minimizing the free energy described in Equations 5-9?” The entity minimizing free energy is not the individual but rather the collective agents who share the variable $w$ defined in Equation 5, representing a collective form of life that engages in predictive coding of the world. While System 3’s operation corresponds to the dynamics of symbol and language emergence, one might assume that its learning relies solely on System 2 processes, such as individual language learning. However, we assert that System 3 is a collective emergent system. It adapts to its environment not only through the internal neural plasticity of individual agents but also by leveraging the plasticity of language itself in predictive coding. Consequently, the language formed through this process reflects collective predictive coding as a group.
@@ 3.3 Relationships and Dynamics Among Systems: A Cognitive and Temporal Perspective
4 Extended Bergsonian MTS for System 0/1/2/3
Bergson’s philosophy provides a grounding framework for the above-discussed extended four-tiered system in philosophical terms. This is because his philosophy of time encompasses not only the “fast and slow” processes of conscious cognition addressed in his second major work Matter and Memory, but also issues of evolution and society, corresponding respectively to his third and fourth major works. In this section, we introduce Bergson’s approach, which combines the “landscape,” an extended notion of space, with the “window of the present,” rendered elastic, as a way to structure and clarify the conceptual foundations behind such a wide range of phenomena.
@@ 4.1 Three Axes of Timescales
Bergson’s philosophy of time offers a broad perspective on diverse phenomena, ranging from matter to society, each operating on different timescales. When distinguishing between different these timescales, there are it is beneficial to differentiate between multiple axes that, which should not be conflated (see Figure 5).
@@
The first axis refers to the multiple temporal layers that structure an individual organism. Centered around the present window—framed by layers 1 and 2 as described earlier—, it comprises a material layer below (layer 0) and a memory layer above (layer 3) (see the MTS structure: on the right side of Figure 5). The second axis represents the timescales required for the formation of adaptive optimization (“landscape” formation or “canalization,” discussed later; see left of Figure 5). Evolution, for instance, involves timescales ranging from thousands to millions of years, while learning involves timescales ranging from weeks to months. The third axis represents the size of the present window (middle of Figure 5). A thicker present window enables real-time access to a richer reservoir of pure memory in layer 3. Strictly speaking, this is not a difference in time “scales,” but it serves as a decisive factor influencing higher-order emergent properties, such as deliberation and verbal ingenuity.
In real-world phenomena, these multiple axes of timescales intertwine and manifest together. Taking the systems proposed in this paper as examples: Both Systems 0 and 1 utilize only a “thin” window of the present, but the timescales of the underlying canalization differ between evolution and learning. System 2 has a “thick” present window, enabling access to memory, and interacts with the learning landscape. System 3 has an “extended” present window and additionally interacts with a collectively formed societal landscape.
@@ 4.2 Two Types of Memory and Thickness of the Present
Bergson’s theory of time explains natural phenomena through the combined effect of two types of “memory.” This distinction is often seen as a precursor to modern concepts of procedural and episodic memory [145, 106]; however, it has a broader scope. One type is “motor memory,” which refers to a series of motor patterns learned through repetition, such as skills and habits in humans. The other type is “pure memory,” which preserves original experiences before they become organized into habitual patterns. For example, through repeated trials of bicycle riding (r1, r2, r3), a motor memory for riding a bicycle (R) is acquired. However, one can also recall r1 as a specific life event on a particular date. Pure memory is what retains these unique instances.
Critically, the organization of motor memory, as an adaptation to the environment, operates under the constraints of utility at the given moment, whereas pure memory provides a temporally open resource—offering a surplus beyond these established habitual patterns. This surplus allows for flexibility and adaptability in novel situations that may arise later. Pure memory, with its flexibility and temporal openness, can be conceptually related to external memory systems utilized in recent machine learning approaches, such as Neural Turing Machines and Transformer architectures equipped with external memories.
Let us now revisit the two types of recognition briefly mentioned in Section 2.4: a fast but limited automatic process and a slower but intentional process, the latter termed “attentive recognition” ([9], chap. 2). Both automatic and attentive recognition rely on motor memory and share the function of facilitating interactions within stereotyped environments. However, automatic recognition, operating with a “thin” window of the present, is confined to immediate motor responses. In contrast, attentive recognition, with its “thick” window, provides access to pure memory. This temporal opening allows active reconstruction of the recognized object as a mental image, and in some cases, even enables re-experiencing scenes from one’s personal past. Such temporal distinctions also likely apply to Systems 1 and 2.
Motor memory is essential for streamlining immediate cognition and action. Without this, complex motor responses would need to be assembled from scratch each time, making hierarchical and intricate actions unattainable. However, if all experiences were solely dedicated to constructing motor memory without leaving any surplus, there would be no additional resources to draw upon in novel situations. The size of the present window acts as a constraint on this real-time improvisation. Thus, these dual memory resources are fundamentally important for intellectual pursuits and creativity. They enable preparedness and support through motor memory while retaining the freedom to deviate when necessary.
@@ 4.3 Extended Definition of Motor Memory: Canalized Landscape
In this paper, we define motor memory as self-organized “prior optimization” that occurs before actual experience (here, optimization does not necessarily aim for the global optimum). Mastering a movement and learning its patterns reduces cognitive and motor control demands when a similar situation arises again. Motor memory effectively serves as a form of temporal outsourcing of optimization. This indicates that motor memory is not limited to internal mechanisms, such as updating motor programs or world models within an organism. Rather, organisms actively modify their environment through behaviors such as marking territory or building nests—and in humans, through artifacts like roads, buildings, and product design—to create conditions facilitating automatic responses that do not require conscious deliberation.
@@
This definition, which focuses on the underlying temporal structure rather than the phenomenological classification, also captures adaptive optimization on an evolutionary scale, viewed as a large-scale self-organizing landscape.Such mechanisms lead to unconscious behaviors based on innate behavioral and morphological traits, corresponding to System 0 (Bergson calls this “instinct”).
@@
Bergson cites the example of the Sphex wasp stinging a caterpillar’s nine nerve centers in a precise order to paralyze but not kill it ([10], pp. 174–175). This intricate maneuver is neither learned nor reasoned but pre-configured. On an evolutionary scale, this has been achieved through countless variations and selections. Thus, behavioral optimization is completed even prior to an organism’s birth.
@@
At both individual and evolutionary levels, establishing stable sensorimotor pathways in advance eliminates the need for a present window. In this sense, Bergson argues that as motor memory becomes perfected, it “steps outside of time”([9], p. 91). To illustrate this point further, we introduce the concept of the “Canalized Landscape.”
@@
Bergson’s theory of perception involves the idea of knowledge outsourcing. In Matter and Memory, he rejects both representationalism and atomistic associationism, proposing a form of “direct realism” grounded in an organism’s potential actions. The idea that potential behavioral inducements are structured within the environment not only anticipates von Uexküll’s notion of the Umwelt [148] but also aligns remarkably well with the dynamical systems perspective discussed in Section 2.1. Each organism operates within its own action-inducing space, which extends beyond its body and is shaped through motor memory processes, including evolution and learning. Bergson likened its formation process to river formation over geological timescales, calling it “canalization” ([9], p. 280, [10], pp. 94-95). Following C. H. Waddington, a renowned geneticist who inherited this idea via A. N. Whitehead [107], we term the resultant space the “landscape” [149].
@@
Thanks to this canalized landscape, organisms do not need to build behavior anew each time. Instead, they pre-construct an appropriate landscape in which the potential for appropriate actions is well arranged. Canalization shapes this potential space, and physical laws, evolutionary adaptations, and learning all contribute to it. Ideally, organisms achieve locally optimized behaviors by simply “rolling down” this landscape.
@@
If a river meanders suitably, water does not need to “think”; by flowing towards lower potential, it naturally follows the optimized “valley” paths. Similarly, the Sphex wasp’s behavior arises not solely from internal motor control but also from coupling with the caterpillar’s body shape and structure, allowing its actions to unfold automatically. Here, the spatial distinction between the internal and external becomes less significant. Instead, the focus is the temporally outsourcing of landscape formation (canalization) prior to the window of the present window.